# Kitashirakawa Oiwakecho, Sakyo-ku, Kyoto 606-8502, Japan

### v1

@@ Abstract
In the AdS/CFT correspondence, direct scattering in the bulk may not have a local boundary analog. A nonlocal implementation on the boundary requires $O(1/G_{N})$ mutual information. This statement is formalized by the connected wedge theorem, which can be proven using general relativity within AdS3 but also by applying quantum information theory on the boundary, suggesting that the theorem applies to any holographic duality. We examine scattering within the static patch of asymptotically dS3 spacetime, which is conjectured to be described by a quantum theory on the stretched horizon in static patch holography. We prove that causality on the horizon should be induced from null infinities $\mathcal{I}^{\pm}$ to maintain consistency with the theorem. Specifically, signals propagating in the static patch are associated with local operators at $\mathcal{I}^{\pm}$. Our results suggest a novel connection between static patch holography and the dS/CFT correspondence.
@@ 1 Introduction
The quantum mechanical description of the gravitational aspects of our universe, encompassing its geometry and causality, remains one of the most profound questions in theoretical physics. A lot of progress has been made via the explicit realization of the AdS/CFT correspondence Maldacena:1997re ; Witten:1998qj , where anti-de Sitter (AdS) spacetime is dual to a conformal field theory (CFT) living on its asymptotic boundary. Toward a quantum mechanical description of our universe, which is approximately de Sitter (dS) spacetime, dS holography stands out as a promising but enigmatic framework Strominger:2001pn ; Bousso:2001mw ; Abdalla:2002hg ; Alishahiha:2004md ; Parikh:2004wh ; Alishahiha:2005dj ; McFadden:2009fg ; Dong:2010pm ; Anninos:2011ui ; Anninos:2011af ; Dong:2018cuv ; Gorbenko:2018oov ; Arias:2019pzy ; Arias:2019zug ; Lewkowycz:2019xse ; Geng:2021wcq ; Susskind:2021esx ; Shaghoulian:2021cef ; Susskind:2021omt ; Coleman:2021nor ; Susskind:2021dfc ; Hikida:2022ltr ; Svesko:2022txo ; Levine:2022wos ; Shaghoulian:2022fop ; Lin:2022nss ; Susskind:2022dfz ; Susskind:2022bia ; Banihashemi:2022htw ; Rahman:2022jsf ; Goel:2023svz ; Narovlansky:2023lfz ; Susskind:2023rxm ; Giveon:2023rsk ; Franken:2023pni ; Kawamoto:2023nki ; Galante:2023uyf . One of its fundamental puzzles arises from the fact that dS spacetime is a closed universe, lacking a boundary in the traditional sense. This leads to the question: Where do the holographic degrees of freedom reside?
@@
A natural extension of AdS/CFT to de Sitter spacetime is known as the dS/CFT correspondence Strominger:2001pn ; Bousso:2001mw ; Balasubramanian:2001nb ; Anninos:2011ui , which is achieved via analytical continuation. This approach situates the dual CFT at future or past null infinity, where conventional notions of time and states are ill-defined. Consequently, standard concepts such as entanglement entropy, which are pivotal in understanding quantum gravity Maldacena:2001kr ; Ryu:2006bv ; Hubeny:2007xt ; VanRaamsdonk:2009ar ; VanRaamsdonk:2010pw ; Maldacena:2013xja ; Dong:2016eik ; Penington:2019npb ; Almheiri:2019psf ; Almheiri:2019hni ; Doi:2023zaf , become ambiguous. In this scenario, the “EPR” aspect of the ER=EPR conjecture Maldacena:2013xja is notably absent, complicating the interpretation of quantum entanglement in this context.
@@
An alternative proposal is static patch holography Susskind:2021omt ; Susskind:2021dfc ; Susskind:2021esx ; Shaghoulian:2021cef ; Shaghoulian:2022fop ; Lin:2022nss ; Susskind:2023hnj , which retains a notion of time and unitarity. Here, a holographic screen is situated in the bulk, close to the cosmological horizon of an observer, and encodes the state of the static patch of this observer. The precise nature of the dual theory remains unsettled. While the double-scaled Sachdev-Ye-Kitaev (DSSYK) model is one conjecture Susskind:2021esx ; Lin:2022nss ; Rahman:2022jsf ; Goel:2023svz ; Narovlansky:2023lfz ; Verlinde:2024znh ; Verlinde:2024zrh ; Blommaert:2023opb ; Blommaert:2023wad ; Rahman:2024iiu , it involves dimensional reduction, leaving the description of $d$-dimensional dS spacetimes for $d\geq 3$ unresolved. In particular, the dual theory in a higher-dimensional de Sitter spacetime is expected to be highly nonlocal, a feature that low-dimensional holography escapes. The ambiguity of the spatial extent of the holographic screen further complicates identifying the correct prescription for the holographic entanglement entropy, leading to various competing proposals Susskind:2021esx ; Shaghoulian:2021cef ; Shaghoulian:2022fop ; Franken:2023pni .
@@
To understand a holographic spacetime from a local, ultraviolet (UV) dual quantum theory, identifying the correct location of the UV boundary/screen is inevitable. In static patch holography, a holographic screen is placed away from the asymptotic boundary. Generally, a theory on such a screen is expected to be nonlocal so causality must be treated carefully. Furthermore, the precise understanding of causality on the boundary/screen consistent with the bulk causality is essential to resolve various puzzles related to holographic entanglement entropy such as a violation of subadditivity Kawamoto:2023nki ; Mori:2023swn ; Grado-White:2020wlb ; Lewkowycz:2019xse .
@@
In this spirit, one interesting question is how a holographic quantum theory encodes the causal structure of the bulk dual May:2019yxi ; May:2019odp ; May:2021nrl ; May:2022clu ; Caminiti:2024ctd . Let us consider a situation where we send information from a set of input points, process it, and share the outcome among output points. This procedure, known as a send-receive quantum task, can have a local bulk realization but not necessarily so on the boundary. Whether the task can be performed locally in some background is purely a causality statement so it is entirely answered from the causal structure and locations of input and output points. The connected wedge theorem May:2019odp in AdS3/CFT2 relates this causality statement to correlation, namely, that when a $2$-to-$2$ scattering from input and output points on the boundary is possible in the bulk, but not on the boundary, there must be $O(1/G_{N})$ mutual information between certain boundary causal diamonds, which are fixed by the input and output locations and the boundary causality.
@@
This paper is organized as follows. For readers only interested in our main claim, skip to Section 2.3 for a summary of the main results or Section 5 for further details. In Section 3, we provide a short recap on the connected wedge theorem in asymptotically AdS and one with branes or cutoff surfaces Mori:2023swn . Section 4 briefly reviews the geometry of dS and static patch holography. In Section 5, we present a puzzle where the connected wedge theorem is violated in static patch holography. We then resolve this problem by introducing the notion of an induced lightcone, which is constructed from a point at null infinity. Additionally, we examine how induced causality changes as we push the holographic screen deep inside the static patch. In Section 6, we first show how induced causality resolves the apparent violation of the connected wedge theorem by explicitly calculating holographic entanglement entropy. We then prove the theorem in the static patch of asymptotically dS3 spacetime using induced causality on the holographic screen. Possible future directions are discussed in Section 7.
In Appendix A, we discuss formal aspects related to covariant holographic entanglement entropy prescriptions in static patch holography. In particular, we review the definitions of the extremization and maximin procedure as well as the constrained extremization proposed in Franken:2023pni , and prove their equivalence in the static patch.
@@ 2 Main results and notations
In this section, we list the notation of symbols we use throughout the paper, briefly review the assumptions about the spacetimes, and present a concise summary of our main results.
2.1 Summary of notations
This paper involves several notations. Basically,
  • Plain capital letters $B,D,J$,… denote bulk subregions.
  • The lower case Greek letter $\gamma$ denotes achronal codimension-two surfaces.
  • Lower case letters $c,r,p$,… denote points in the bulk.
  • Calligraphic capital letters $\mathscr{A,L,N,R,S}$,… are used to denote subsystems of the null membrane, an object used in the proof of Theorem 6.1.
  • $X_{M}$ with $M=0,1,2,3$ denotes the embedding coordinates of dS3.
    $\theta=\frac{1}{\mathcal{A}}\frac{d\mathcal{A}}{d\lambda},$ (1)
  • $\mathcal{I}^{\pm}$ are the past ()(-)( - ) and future (+)(+)( + ) null infinities of $dS$. Points or regions defined on $\mathcal{I}^{\pm}$ are tilted with ${}\tilde{~}{}$.
@@ 2.3 Brief summary of main results
In this paper, we present a puzzle in pure dS3 where the connected wedge theorem is violated. We consider a $2$-to-$2$ scattering with input and output points on the cosmological horizon of an observer, where the holographic screen is supposed to be located by the static patch holography proposal. [^2] Sometimes we place a Planck-distance cutoff, replacing the screen with the stretched horizon. Even in such a case, the same puzzle occurs and the general proof presented in the later section will apply to both cases. For simplicity, we choose these points such that the scattering in the bulk barely occurs – the scattering region is pointlike. By tuning the spatial location of the input and output points, one can create a situation where the $2$-to-$2$ scattering is impossible on the screen, at least from local interactions. The connected wedge theorem predicts that a quantum task via scattering is non-locally implemented on the screen by $O(1/G_{N})$ entanglement between two causally disconnected screen subsystems. However, the highly constrained causality due to the lightlike nature of the holographic screen leads us to conclude that these subsystems are too small to accommodate the expected large entanglement based on the holographic entanglement entropy proposal for static patch holography Shaghoulian:2021cef ; Franken:2023pni .
@@
To resolve this puzzle, we need to understand which screen causality is being used. We distinguish three different causalities on the screen. 1) Causality based on the induced metric of the screen (which will be denoted by a subscript S), 2) causality determined from the intersection of the bulk lightcone of a point on the screen and the screen itself (which will be denoted by a conditioning $(\cdot)|_{\mathcal{S}}$), and 3) induced causality, defined as an intersection between the bulk lightcone from a null infinity and the screen (which will be denoted by an accent $\hat{\phantom{a}}$). The first type and the second type of causalities may look similar, however, they are different in general. In Section 5, we present a case where they are indeed different.
@@
Inspired by a recent work by one of the authors Mori:2023swn , we propose the consistent resolution of the puzzle is given by the causality induced by a local operator from the null infinities. See Figure 1 for a visualization. We then show the induced causality as defined in Section 5 resolves the apparent violation of the connected wedge theorem as the induced lightcone (which should not be confused with a lightcone constructed from the induced metric) extends farther than the local lightcones, leading to larger causal regions associated with the scattering. This leads us to state a general version of the connected wedge theorem in the static patch of three-dimensional asymptotically de Sitter spacetime. Before stating the theorem, let us first define the induced causality:
@@
where the minimization is taken so that the minimized tilted point $\tilde{c}=\tilde{c}_{\ast}$ satisfies $\hat{J}^{+}(\tilde{c}_{\ast})\subseteq\hat{J}^{+}(\tilde{c}^{\prime})$ for any $\tilde{c}^{\prime}$. Analogously, we define a past induced lightcone by replacing $c$ with $r$ and $+$ with --.
@@
Let $\mathcal{S}$ be the holographic screen in the static patch of an asymptotically dS3 spacetime. Assuming static patch holography and its associated entropy prescription, if the $2$-to-$2$ scattering $c_{1},c_{2}\in\mathcal{S}\rightarrow r_{1},r_{2}\in\mathcal{S}$ is possible in the bulk but not on the screen $\mathcal{S}$ based on the induced causality, then $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ have mutual information as large as $O(1/G_{N})$, implying their entanglement wedge is connected in the bulk bounded by $\mathcal{S}$.
A precise version of this theorem is given in Theorem 6.1. This statement is proven using the notion of induced causality on the holographic screen. The proof is analogous to that of May:2019odp ; Mori:2023swn , using the extremality and the equivalence to the maximin prescription.
@@
Finally, the appendix of this paper contains formal results regarding the subtleties in the definition of holographic entropy prescription(s) in static patch holography. Identifying the prescription is of the highest importance in the context of the connected wedge theorem, as the proof relies on the extremality of the surface computing entropy as well as the equivalence of the surface to the maximin one. However, this is not trivial in the proposed dS holographic entropy prescriptions. So far two proposals have been made: the monolayer and bilayer proposals Susskind:2021esx ; Shaghoulian:2021cef ; Shaghoulian:2022fop . The monolayer proposal appears inconsistent with the entanglement wedge reconstruction as pointed out in a paper by one of the authors Franken:2023pni . However, even if we adopt the bilayer proposal, there remains an ambiguity regarding the ‘extremization’. In contrast to AdS/CFT, where an extremized surface always exists and is equivalent to the maximin surface Wall:2012uf , the extremization problem may have no solution in static patch holography due to the non-asymptotic feature of the holographic screen. This can be resolved by relaxing the global extremizing condition, leading to the so-called constrained extremization Franken:2023pni . However, there remains a subtlety as it is not guaranteed equivalent to the maximin surface as in AdS/CFT. Indeed, we prove that they are different in general (Theorem A.3). Nevertheless, we find these three definitions (extremization, constrained extremization, and maximin) are equivalent within the static patch (Theorem A.5).

### v3

@@ Abstract
In the AdS/CFT correspondence, a direct scattering in the bulk may not have a local boundary analog. A nonlocal implementation on the boundary requires $O(1/G_{N})$ mutual information. This statement is formalized by the connected wedge theorem, which can be proven using general relativity within AdS3 but also argued for using quantum information theory on the boundary, suggesting that the theorem applies to any holographic duality. We examine scattering within the static patch of asymptotically dS3 spacetime, which is conjectured to be described by a quantum theory on the stretched horizon in static patch holography. We show that causality on the horizon induced from null infinities $\mathcal{I}^{\pm}$ is consistent with the theorem. Specifically, signals propagating in the static patch are associated with local operators at $\mathcal{I}^{\pm}$. Our results suggest a novel connection between static patch holography and the dS/CFT correspondence.
@@ 1 Introduction
The quantum mechanical description of the gravitational aspects of our universe, encompassing its geometry and causality, remains one of the most profound questions in theoretical physics. A lot of progress has been made via the explicit realization of the AdS/CFT correspondence Maldacena:1997re ; Witten:1998qj , where anti-de Sitter (AdS) spacetime is dual to a conformal field theory (CFT) living on its asymptotic boundary. In parallel to these developments, there has been strong evidence that our universe has gone through two periods of accelerated expansion, which are the cosmic inflation in the early universe and the present period SupernovaSearchTeam:1998fmf ; SupernovaCosmologyProject:1998vns ; Planck:2015fie . An extensive understanding of the early universe asks for a quantum gravitational description, and extending the great progress made in AdS/CFT to expanding universes is one of the major challenges in modern high-energy physics. Toward a quantum mechanical description of our universe, which is approximately de Sitter (dS) spacetime, dS holography stands out as a promising but enigmatic framework Strominger:2001pn ; Bousso:2001mw ; Abdalla:2002hg ; Alishahiha:2004md ; Parikh:2004wh ; Alishahiha:2005dj ; McFadden:2009fg ; Dong:2010pm ; Anninos:2011ui ; Anninos:2011af ; Dong:2018cuv ; Gorbenko:2018oov ; Arias:2019pzy ; Arias:2019zug ; Lewkowycz:2019xse ; Geng:2021wcq ; Susskind:2021esx ; Shaghoulian:2021cef ; Susskind:2021omt ; Coleman:2021nor ; Susskind:2021dfc ; Hikida:2022ltr ; Svesko:2022txo ; Levine:2022wos ; Shaghoulian:2022fop ; Lin:2022nss ; Susskind:2022dfz ; Susskind:2022bia ; Banihashemi:2022htw ; Rahman:2022jsf ; Goel:2023svz ; Narovlansky:2023lfz ; Susskind:2023rxm ; Giveon:2023rsk ; Franken:2023pni ; Kawamoto:2023nki ; Galante:2023uyf . One of its fundamental puzzles arises from the fact that dS spacetime is a closed universe, lacking a boundary in the traditional sense. This leads to the question: Where do the holographic degrees of freedom reside?
@@
A natural extension of AdS/CFT to de Sitter spacetime is known as the dS/CFT correspondence Strominger:2001pn ; Bousso:2001mw ; Balasubramanian:2001nb ; Anninos:2011ui , which is achieved via analytical continuation. This approach situates the dual CFT at future or past null infinity, where conventional notions of time and states are ill-defined. Consequently, standard concepts such as entanglement entropy, which are pivotal in understanding quantum gravity Maldacena:2001kr ; Ryu:2006bv ; Hubeny:2007xt ; VanRaamsdonk:2009ar ; VanRaamsdonk:2010pw ; Maldacena:2013xja ; Dong:2016eik ; Penington:2019npb ; Almheiri:2019psf ; Almheiri:2019hni ; Doi:2023zaf , become ambiguous.
@@
An alternative proposal is static patch holography Susskind:2021omt ; Susskind:2021dfc ; Susskind:2021esx ; Shaghoulian:2021cef ; Shaghoulian:2022fop ; Lin:2022nss ; Susskind:2023hnj , in which the dual quantum theory is defined on a holographic screen, namely, a codimension-one timelike hypersurface. Under this proposal, the notion of time and unitarity is manifest. Here, a holographic screen is situated in the bulk, close to the cosmological horizon of an observer, and encodes the state of the static patch of this observer. The precise nature of the dual theory remains unsettled. While the double-scaled Sachdev-Ye-Kitaev (DSSYK) model is one conjecture Susskind:2021esx ; Lin:2022nss ; Rahman:2022jsf ; Goel:2023svz ; Narovlansky:2023lfz ; Verlinde:2024znh ; Verlinde:2024zrh ; Blommaert:2023opb ; Blommaert:2023wad ; Rahman:2024iiu ; Milekhin:2023bjv ; Xu:2024hoc ; Milekhin:2024vbb , it involves dimensional reduction, leaving the description of $d$-dimensional dS spacetimes for $d\geq 3$ unresolved. In particular, the dual theory in a higher-dimensional de Sitter spacetime is expected to be highly nonlocal, a feature that low-dimensional holography escapes. The ambiguity of the spatial extent of the holographic screen further complicates identifying the correct prescription for the holographic entanglement entropy, leading to various competing proposals Susskind:2021esx ; Shaghoulian:2021cef ; Shaghoulian:2022fop ; Franken:2023pni ; Hao:2024nhd .
@@
To understand a holographic spacetime from a ultraviolet (UV) dual quantum theory, a natural approach is to identify a location for the dual degrees of freedom to live. In static patch holography, a holographic screen is placed away from the asymptotic boundary. Generally, a theory on such a screen is expected to be nonlocal so causality must be treated carefully. Furthermore, the precise understanding of causality on the boundary/screen consistent with the bulk causality is essential to resolve various puzzles related to holographic entanglement entropy such as a violation of subadditivity Kawamoto:2023nki ; Mori:2023swn ; Grado-White:2020wlb ; Lewkowycz:2019xse .
@@
In this spirit, one interesting question is how a holographic quantum theory encodes the causal structure of the bulk dual May:2019yxi ; May:2019odp ; May:2021nrl ; May:2022clu ; Caminiti:2024ctd . Let us consider a situation where we send information from a set of input points, process it, and share the outcome among output points. This procedure, known as a relativistic quantum task, can have a local bulk realization but not necessarily on the boundary. Whether the task can be performed locally in some background is purely a causality statement so it is entirely answered from the causal structure and locations of input and output points. The connected wedge theorem May:2019odp in AdS3/CFT2 relates this causality statement to correlation, namely, that when a $2$-to-$2$ scattering from input and output points on the boundary is possible in the bulk, but not on the boundary, there must be $O(1/G_{N})$ mutual information between certain boundary causal diamonds, which are fixed by the input and output locations and the boundary causality.
@@
This paper is organized as follows. For readers only interested in our main claim, skip to Section 2.2 for a summary of the main results or Section 5 for further details. In Section 2, we provide the assumptions we make about the spacetimes in the paper and review our main results. In Section 3, we provide a short recap on the connected wedge theorem in asymptotically AdS and one with branes or cutoff surfaces Mori:2023swn . Section 4 briefly reviews the geometry of dS and static patch holography. In Section 5, we present a puzzle where the connected wedge theorem is violated in static patch holography. We then resolve this problem by introducing the notion of an induced lightcone, which is constructed from a point at null infinity. Additionally, we examine how induced causality changes as we push the holographic screen deep inside the static patch. In Section 6, we first show how induced causality resolves the apparent violation of the connected wedge theorem by explicitly calculating holographic entanglement entropy. We then prove the theorem in the static patch of asymptotically dS3 spacetime using induced causality on the holographic screen. Possible future directions are discussed in Section 7.
In Appendix A, we list the notation of symbols we use throughout the paper. In Appendix B, we discuss formal aspects related to covariant holographic entanglement entropy prescriptions in static patch holography. In particular, we review the definitions of the extremization and maximin procedure as well as the constrained extremization proposed in Franken:2023pni , and prove their equivalence in the static patch.
@@ 2 Main results and notations
In this section, we briefly review the assumptions about the spacetimes, and present a concise summary of our main results.
@@ 2.1 Assumptions about the spacetimes
Later we will introduce a holographic screen near the cosmological horizon to discuss static patch holography. Since our scope in this paper is scattering by probe particles, we fix the background geometry and place a screen at a fixed location. Thus, in this paper we are agnostic about the precise boundary condition imposed on the screen although if we consider any perturbation, we need to identify a sensible boundary condition with stability, diffeomorphism, etc. We refer to Anninos:2024wpy for a study on boundary conditions in static patch and Karch:2013oqa for a possible boundary condition on the screen.
@@ 2.2 Brief summary of main results
In this paper, we present a puzzle in pure dS3 where the connected wedge theorem is violated. We consider a $2$-to-$2$ scattering with input and output points on the cosmological horizon of an observer, where the holographic screen is supposed to be located by the static patch holography proposal. [^1] Sometimes we place a Planck-distance cutoff, replacing the screen with the stretched horizon. Even in such a case, the same puzzle occurs and the general proof presented in the later section will apply to both cases. For simplicity, we choose these points such that the scattering in the bulk barely occurs – the scattering region is pointlike. See Figure 1 for a schematic picture.
Figure 1: A schematic picture of a 2-to-2 scattering in the static patch of pure dS3, from input points $c_{1},c_{2}$ to output points $r_{1},r_{2}$ on the cosmological horizon. The white box denoted $U$ corresponds to a unitary operation acting on the signal qubits in the bulk. In both pictures, $\sigma$ indicates the time direction. (a) Picture of the scattering in the Penrose diagram of de Sitter space. Each slice of the diagram has the topology of a sphere and the vertical lines on the edges follow the trajectories of static observers on the poles of the sphere. The diagonal lines are the cosmological horizons of the two observers. (b) Picture of the scattering in the static patch. Each cosmological horizon (or stretched horizon) has the topology of a cylinder and the bulk (static patch) fills inside. Input and output points are chosen such that a direct scattering is only possible at one bulk point. The angular coordinate on the cylinder is denoted by $\varphi$.
By tuning the spatial location of the input and output points, one can create a situation where the $2$-to-$2$ scattering is impossible on the screen, at least from local interactions. The connected wedge theorem predicts that a quantum task via scattering is non-locally implemented on the screen by $O(1/G_{N})$ entanglement between two causally disconnected screen subsystems. However, the highly constrained causality due to the lightlike nature of the holographic screen leads us to conclude that these subsystems are too small to accommodate the expected large entanglement based on the holographic entanglement entropy proposal for static patch holography Shaghoulian:2021cef ; Franken:2023pni .
@@
To resolve this puzzle, we need to understand which screen causality is being used. We distinguish three different causalities on the screen. 1) Causality based on the induced metric of the screen $\mathcal{S}$ (which will be denoted by a subscript S), 2) causality determined from the intersection of the bulk lightcone of a point on the screen and the screen itself (which will be denoted by conditioning $(\cdot)|_{\mathcal{S}}$), and 3) induced causality, defined as an intersection between the bulk lightcone from a null infinity and the screen (which will be denoted by an accent $\hat{\phantom{a}}$). The first type and the second type of causalities may look similar, however, they are different in general. In Section 5, we present a case where they are indeed different.
@@
Inspired by a recent work by one of the authors Mori:2023swn , we propose the consistent resolution of the puzzle is given by the causality induced by a local operator from the null infinities. We then show the induced causality as defined in Section 5 resolves the apparent violation of the connected wedge theorem as the induced lightcone (which should not be confused with a lightcone constructed from the induced metric) extends farther than the local lightcones, leading to larger causal regions associated with the scattering. This leads us to state a general version of the connected wedge theorem in the static patch of three-dimensional asymptotically de Sitter spacetime. [^2] In this paper, we often use the term static patch to describe the causal region of an observer, by analogy with the case of the pure de Sitter spacetime. Thus, the geometry of this bulk region is not necessarily described by a static metric. Before stating the theorem, let us first define the induced causality:
@@
where the minimization is taken so that the minimizing point $\tilde{c}=\tilde{c}_{\ast}$ satisfies $J^{+}(\tilde{c}_{\ast})\cap\mathcal{S}\subseteq J^{+}(\tilde{c}^{\prime})\cap\mathcal{S}$ for any $\tilde{c}^{\prime}$. Analogously, we define a past induced lightcone by replacing $c$ with $r$ and $+$ with --. See Figure 2 for a visualization.
The existence of the minimizing points $\tilde{c},\tilde{r}$ is guaranteed from the definition of null infinities $\mathcal{I}^{\pm}$. Now, let us state the connected wedge theorem in asymptotically dS3:
@@
Let $\mathcal{S}$ be the holographic screen in the static patch of an asymptotically dS3 spacetime. Assuming static patch holography and its associated entropy prescription, if the $2$-to-$2$ scattering $c_{1},c_{2}\in\mathcal{S}\rightarrow r_{1},r_{2}\in\mathcal{S}$ is possible in the bulk but not on the screen $\mathcal{S}$ based on the induced causality, then the decision regions $\mathcal{R}_{1}{\equiv\hat{J}^{+}(c_{1})\cap\hat{J}^{-}(r_{1})\cap\hat{J}^{-}(r_{2})}$ and $\mathcal{R}_{2}{\equiv\hat{J}^{+}(c_{2})\cap\hat{J}^{-}(r_{1})\cap\hat{J}^{-}(r_{2})}$ have mutual information as large as $O(1/G_{N})$, implying their entanglement wedge is connected in the bulk bounded by $\mathcal{S}$.
A precise version of this theorem is given in Theorem 6.1. This statement is proven using the notion of induced causality on the holographic screen. The proof is analogous to that of May:2019odp ; Mori:2023swn , using the extremality and its equivalence to the maximin prescription. Note that our proof of the connected wedge theorem applies not only to scattering among points on the cosmological horizon but also on a general holographic screen, defined in Definition 4.1 and Definition 4.2.
@@
Finally, the appendix of this paper contains formal results regarding the subtleties in the definition of holographic entropy prescription(s) in static patch holography. Identifying the prescription is of the highest importance in the context of the connected wedge theorem, as the proof relies on the extremality of the surface computing entropy as well as the equivalence of the surface to the maximin one. However, this is not trivial in the proposed dS holographic entropy prescriptions. So far two proposals have been made: the monolayer and bilayer proposals Susskind:2021esx ; Shaghoulian:2021cef ; Shaghoulian:2022fop . The monolayer proposal appears inconsistent with the entanglement wedge reconstruction as pointed out in a paper by one of the authors Franken:2023pni . However, even if we adopt the bilayer proposal, there remains an ambiguity regarding the ‘extremization’. In contrast to AdS/CFT, where an extremized surface always exists and is equivalent to the maximin surface Wall:2012uf , the extremization problem may have no solution in static patch holography due to the non-asymptotic feature of the holographic screen. This can be resolved by relaxing the global extremizing condition, leading to the so-called constrained extremization Franken:2023pni . However, there remains a subtlety as it is not guaranteed equivalent to the maximin surface as in AdS/CFT. Indeed, we prove that they are different in general (Theorem B.3). Nevertheless, we find these three definitions (extremization, constrained extremization, and maximin) are equivalent within the static patch (Theorem B.5).